OpenStudy (anonymous):

Which of the following is the equation of a circle whose center is at the origin and whose radius is 4? x2+ y2= 2 x2+ y2= 4 x2+ y2= 16

ganeshie8 (ganeshie8):

equation of circle wid radius \(r\) : \(x^2 + y^2 = r^2\)

OpenStudy (whpalmer4):

formula for a circle with center at \((h,k)\) and radius \(r\) is \[(x-h)^2+(y-k)^2=r^2\] Your center is the origin \((0,0)\) so it reduces to @ganeshie8's form